# Minimal criteria for continuous-variable genuine multipartite entanglement

### v1

@@ IV.1 Left-hand side of the criterion
[…]
where $\ell^{\alpha}$ is a lower triangular matrix with non-zero entries and strictly positive diagonal elements. Ascertaining of the structure of the adjacency matrix $A$, which leads to the quadrature combinations (21) with at most two non-zero terms, is now simple. The quadrature combinations (21) then contain only $24N-$ 2 elements. From the formula (29) it follows immediately that this happens if the vectors $A_{i+1:N,i}$, $i=1,2,\ldots,N-1$, contain only one non-zero element. As the structure of the adjacency matrix is for a given tree dictated by the labeling of its vertices, the question of the possibility to have a GME criterion based only on at most two-mode quadrature combinations boils down to the question of the existence of the tree labeling which would possess in each vector $A_{i+1:N,i}$ only one non-zero component. In what follows we show that such a labeling really exists, at least for some trees.
IV.1.1 Linear tree
@@
$U^{\alpha}_{\rm lin}=\sum_{i=1}^{N-1}\langle[\Delta(\ell^{\alpha}_{ii}\alpha_{i}+\ell^{\alpha}_{i+1i}\alpha_{i+1})]^{2}\rangle+(\ell^{\alpha}_{NN})^{2}\langle(\Delta\alpha_{N})^{2}\rangle,$
@@
IV.1.2 Star tree
Let us now move to the $N$-vertex star tree. We see that the labeling we are looking for is established if we label the central vertex with $N$ and the other vertices are labeled by the natural numbers $N-1,N-2,\ldots,1$ in a decreasing order in the anticlockwise direction (see labeled trees in the third, fifth and eight row of Tab. 3 for examples corresponding to $N=4,5$ and $6$). For this labeling the only non-zero elements of the adjacency matrix are $A_{Ni}=A_{iN}=1$, and they yield sums (22) of the following form:
$U^{\alpha}_{\rm star}=\sum_{i=1}^{N-1}\langle[\Delta(\ell^{\alpha}_{ii}\alpha_{i}+\ell^{\alpha}_{Ni}\alpha_{N})]^{2}\rangle+(\ell^{\alpha}_{NN})^{2}\langle(\Delta\alpha_{N})^{2}\rangle,$
@@ IV.1.3 Generic tree
 $N$ Unlabeled tree  Type of center Rooted tree
$7$
@@
IV.2 Right-hand side of the criterion
@@ IV.2.1 Product criterion
$U^{x}U^{p}$ $\stackrel{{\scriptstyle 1}}{{\geq}}$ $\left(\sum_{k=1}^{K_{N}}\lambda_{k}U_{k}^{x}\right)\left(\sum_{l=1}^{K_{N}}\lambda_{l}U_{l}^{p}\right)$ (35)
$\stackrel{{\scriptstyle 2}}{{\geq}}$ $\left(\sum_{k=1}^{K_{N}}\lambda_{k}\sqrt{U_{k}^{x}U_{k}^{p}}\right)^{2},$
@@
$\stackrel{{\scriptstyle 3}}{{\geq}}\left\{\sum_{l}p_{l}^{(k)}\left[\sqrt{\langle(\Delta u_{i,I_{k}}^{x})^{2}\rangle_{\rho_{I_{k}}^{(l)}}\langle(\Delta u_{j,I_{k}}^{p})^{2}\rangle_{\rho_{I_{k}}^{(l)}}}\right.\right.$
@@
=$\stackrel{{\scriptstyle 5}}{{=}}\frac{1}{4}\left\{\left[(L^{x})^{T}L^{p}\right]_{ij}^{2}+2(|X_{ij}^{(k)}|-X_{ij}^{(k)})\right\},$ (40)
@@
Combining now Eq. (37) with inequality (IV.2.1) we get immediately the following inequality
@@
Finally, putting inequalities (35) and (44) together, we arrive at the sought necessary condition for biseparability of the form:
@@ IV.2.2 Sum criterion
=$\stackrel{{\scriptstyle 2}}{{=}}\sum_{\alpha=x,p}\sum_{l}p_{l}^{(k)}\left[\langle(\Delta u_{i,I_{k}}^{\alpha})^{2}\rangle_{\rho_{I_{k}}^{(l)}}+\langle(\Delta u_{i,J_{k}}^{\alpha})^{2}\rangle_{\rho_{J_{k}}^{(l)}}\right]$
@@
where inequality $1$ again follows from concavity of the variance, whereas in equality $2$ we used the decomposition (38) and the fact that we average over the product state $\rho_{I_{k}\otimes J_{k}}^{(l)}$. To get inequality $3$ we used the inequality between the arithmetic and geometric mean, the uncertainty relations, and the definitions (41) and (42). Finally, if we combine inequalities (48) and (IV.2.2) we obtain the desired sum criterion in the following form:
@@
IV.3 Direct construction of the sum criterion
@@
IV.3.1 Construction of the left-hand side
The quantities $U^{x,p}$, Eq. (22), on the LHS of both criteria are the same and their construction is easy. Namely, each $N$-vertex tree consists of $N-1$ edges [44]. Each of the edges connects a pair of vertices with labels $i$ and $j$, where by $i$ we denote the label which is less than the label $j$, i.e., $i<j$. With each of the edges we associate a two-mode quadrature linear combination $u_{i}^{\alpha}$ (see Fig. 3). Sum of the variances $\langle(\Delta u_{i}^{\alpha})^{2}\rangle$ over all edges then comprises (up to the term $(\ell^{\alpha}_{NN})^{2}\langle(\Delta\alpha_{N})^{2}\rangle$) the quantity $U^{\alpha}$.
@@ IV.3.1 Construction of the left-hand side
The quantity $U^{\alpha}$ in Eq. (22) is then given by the sum of variances of all $N-1$ quantities (54), plus the term $(\ell^{\alpha}_{NN})^{2}\langle(\Delta\alpha_{N})^{2}\rangle$. Making use of this rule we can easily rederive the quantities $U^{\alpha}$ for the linear tree and the star tree given in Eqs. (IV.1.1) and (IV.1.2), respectively. In the next section we also use the rule to construct the criteria corresponding to the more complex trees in the third column of Tab. 3.
IV.3.2 Construction of the right-hand side
@@
Moving to the construction of the RHS it is obvious that to every bipartite split $k\equiv I_{k}|J_{k}$ of $N$ modes into two disjoint subsets $I_{k}=\{i_{1},i_{2},\ldots,i_{l}\}$ and $J_{k}=\{i_{l+1},i_{l+2},\ldots,i_{N}\}$ there corresponds exactly the same bipartite split of the set of all $N$ vertices of the considered graph. The split divides the set of all $N-1$ pairs $(i,e_{ij})$ into two disjoint subsets, denoted as $\mathcal{U}$ and $\mathcal{C}$, depending on whether the edge $e_{ij}$ is uncut by the split or it is cut off by the split. Formally, $(i,e_{ij})\in\mathcal{U}$ if $i,j\in I_{k}$ or $i,j\in J_{k}$, whereas $(i,e_{ij})\in\mathcal{C}$ if $i\in I_{k}$ and $j\in J_{k}$ or vice versa. Now, to each pair $(i,e_{ij})\in\mathcal{U}$ we ascribe the term $|\ell^{x}_{ii}\ell^{p}_{ii}+\ell^{x}_{ji}\ell^{p}_{ji}|$, whereas with each pair $(i,e_{ij})\in\mathcal{C}$ we associate a different term $|\ell^{x}_{ii}\ell^{p}_{ii}|+|\ell^{x}_{ji}\ell^{p}_{ji}|$. The sum of all the latter terms for all $N-1$ pairs $(i,e_{ij})$ plus the term $|\ell^{x}_{NN}\ell^{p}_{NN}|$, which corresponds to the root, gives finally the quantity $\mathcal{K}^{(k)}$, Eq. (51), appearing on the RHS of the sum criterion (50).
@@
V Applications
V.1 Three-mode states
First, we test functionality of the product criterion (47) and sum criterion (50) for the simplest case of $N=3$ modes. Then there is only one linear tree which for the standard labeling yields the adjacency matrix given above Eq. (IV.1.1). Hence, one gets using the formula (45)
$\mathcal{L}$ $=$ $(\ell^{x}_{11}\ell^{p}_{11}+\ell^{x}_{21}\ell^{p}_{21})^{2}+(\ell^{x}_{21}\ell^{p}_{22})^{2}+(\ell^{x}_{22}\ell^{p}_{21})^{2}$ (55)
@@ V.1 Three-mode states
by Eq. (46). Making use of the expressions for the quantities $U_{\rm lin}^{x,p}$, Eq. (IV.1.1), we get finally the minimal three-mode product criterion for GME,
$\prod_{\alpha=x,p}\left\{\sum_{i=1}^{2}\langle[\Delta(\ell^{\alpha}_{ii}\alpha_{i}+\ell^{\alpha}_{i+1i}\alpha_{i+1})]^{2}\rangle+\left(\ell^{\alpha}_{33}\right)^{2}\langle(\Delta\alpha_{3})^{2}\rangle\right\}$
@@
$\sum_{\alpha=x,p}\left\{\sum_{i=1}^{2}\langle[\Delta(\ell^{\alpha}_{ii}\alpha_{i}+\ell^{\alpha}_{i+1i}\alpha_{i+1})]^{2}\rangle+\left(\ell^{\alpha}_{33}\right)^{2}\langle(\Delta\alpha_{3})^{2}\rangle\right\}$
$\geq\mathrm{min}\left(|\ell^{x}_{11}\ell^{p}_{11}|+|\ell^{x}_{21}\ell^{p}_{21}|+|\ell^{x}_{22}\ell^{p}_{22}+\ell^{x}_{32}\ell^{p}_{32}|+|\ell^{x}_{33}\ell^{p}_{33}|,\right.$
$\left.|\ell^{x}_{11}\ell^{p}_{11}+\ell^{x}_{21}\ell^{p}_{21}|+|\ell^{x}_{22}\ell^{p}_{22}|+|\ell^{x}_{32}\ell^{p}_{32}|+|\ell^{x}_{33}\ell^{p}_{33}|\right).$
At the outset we confirm on a particular example that a fully inseparable but not GME state really does not violate the inequalities (V.1) and (V.1). As an example we use the family of fully inseparable biseparable states (13) with CM $\gamma^{\rm test}$, Eq. (14), where $r\in(0,1.24)$. For all states with the squeezing parameter $r_{i}=0.05\times i$, $i=0,1,\ldots,24$, we have minimized numerically the differences $D_{P}$ and $D_{S}$, Eqs. (52) and (53) with $N=3$, i.e., the difference of the LHS and the RHS of inequalities (V.1) and (V.1). Our analysis reveals that for all considered values of the squeezing the differences are very small but always strictly positive numbers, which confirms correctness of the proposed criteria.
@@ V.1.1 Analytical examples
$\gamma^{\rm sss}=\frac{1}{4}\left[\left(\begin{array}[]{ccc}a_{+}&b_{+}&-b_{+}\\ b_{+}&c_{+}&d_{+}\\ -b_{+}&d_{+}&c_{+}\end{array}\right)\oplus\left(\begin{array}[]{ccc}a_{-}&b_{-}&-b_{-}\\ b_{-}&c_{-}&d_{-}\\ -b_{-}&d_{-}&c_{-}\end{array}\right)\right],$ (59)
@@
Again, we minimized the difference $D_{P}$, Eq. (52), over the parameters $\ell^{\alpha}_{ij}$. To speed up the numerical search we further imposed on the parameters the following constraint $-1\leq\ell^{\alpha}_{ij}\leq 1$. The corresponding minimal differences for the CMs (10) and (59) with squeezing parameters $r_{i}=0.05\times i$, $i=0,1,\ldots,40$ are depicted in Fig. 5. The figure as well as the numerics reveal that for the GHZ-like state except for the case $r=0$ for all other considered values of the squeezing the minimized difference is negative. Similarly, in the case of the split squeezed state the minimized difference becomes negative as soon as the considered value of squeezing exceeds the value of $r=0.6$. These two examples clearly demonstrate the ability of the proposed product criterion (V.1) to detect GME of basic three-mode Gaussian states.
@@
Interestingly, the sum criterion (V.1) fails to detect GME of any of the previous three-mode states detected by the product criterion. Nevertheless, one can find examples of three-mode GME Gaussian states, which are detected by both criteria or only by the sum criterion itself.
V.1.2 Numerical examples
We find examples of Gaussian states whose GME can be detected by the sum criterion (V.1) using the Gaussian analog [35] of the iterative algorithm [33]. Initially, a pure-state CM $\gamma^{(0)}$ is randomly generated. In the first step of the algorithm an optimal witness $Z^{(1)}$ detecting GME from minimal set of two-mode marginal CMs of CM $\gamma^{(0)}$ is found. The second step consists of finding an optimal CM $\gamma^{(1)}$ minimizing $\mbox{Tr}(\gamma Z^{(1)})$. In the next iteration the output CM $\gamma^{(1)}$ is taken as an input to the first step which produces an optimal witness $Z^{(2)}$. The new witness is then used in the second step to get optimal CM $\gamma^{(2)}$ for the witness, etc. By performing few iterations of the algorithm on a randomly generated seed CMs we produced one hundred CMs of GME states, which were then tested by our product and sum GME criteria (V.1) and (V.1). In total, our criteria were capable of detecting GME for nearly 80% of the CMs. One of the obtained CMs is given by

### v3

@@ 4.1 Left-hand side of the criterion
$L^{\alpha}=(\mathbbm{1}_{N}+A)\circ\ell^{\alpha},$ (26)
where $\ell^{\alpha}$ is a lower triangular matrix with non-zero entries and strictly positive diagonal elements. Ascertaining of the structure of the adjacency matrix $A$, which leads to the quadrature combinations (19) with at most two non-zero terms, is now simple. The quadrature combinations (19) then contain only 4N24N-2 elements. From the formula (25) it follows immediately that this happens if the vectors $A_{i+1:N,i}$, $i=1,2,\ldots,N-1$, contain only one non-zero element. As the structure of the adjacency matrix is for a given tree dictated by the labeling of its vertices, the question of the possibility to have a GME criterion based only on at most two-mode quadrature combinations boils down to the question of the existence of the tree labeling which would possess in each vector $A_{i+1:N,i}$ only one non-zero component. In what follows we show that such a labeling really exists, at least for some trees.
4.1.1 Linear tree
@@
$U^{\alpha}_{\rm lin}=$ $\sum_{i=1}^{N-1}\langle[\Delta(\ell^{\alpha}_{ii}\alpha_{i}+\ell^{\alpha}_{i+1i}\alpha_{i+1})]^{2}\rangle$
$+(\ell^{\alpha}_{NN})^{2}\langle(\Delta\alpha_{N})^{2}\rangle,$ (27)
@@
4.1.2 Star tree
Let us now move to the $N$-vertex star tree. We see that the labeling we are looking for is established if we label the central vertex with $N$ and the other vertices are labeled by the natural numbers $N-1,N-2,\ldots,1$ in a decreasing order in the anticlockwise direction (see labeled trees in the third, fifth and eighth row of Tab. 3 for examples corresponding to $N=4,5$ and $6$). For this labeling the only non-zero elements of the adjacency matrix are $A_{Ni}=A_{iN}=1$, and they yield sums (20) of the following form:
$U^{\alpha}_{\rm star}=$ $\sum_{i=1}^{N-1}\langle[\Delta(\ell^{\alpha}_{ii}\alpha_{i}+\ell^{\alpha}_{Ni}\alpha_{N})]^{2}\rangle$
$+(\ell^{\alpha}_{NN})^{2}\langle(\Delta\alpha_{N})^{2}\rangle,$ (28)
@@ 4.1.3 Generic tree
 NN Unlabeled tree  Type of center Rooted tree
77
@@
4.2 Right-hand side of the criterion
@@ 4.2.1 Product criterion
$U^{x}U^{p}$ $\stackrel{{\scriptstyle 1}}{{\geq}}\left(\sum_{k=1}^{K_{N}}\lambda_{k}U_{k}^{x}\right)\left(\sum_{l=1}^{K_{N}}\lambda_{l}U_{l}^{p}\right)$
$\stackrel{{\scriptstyle 2}}{{\geq}}\left(\sum_{k=1}^{K_{N}}\lambda_{k}\sqrt{U_{k}^{x}U_{k}^{p}}\right)^{2},$ (31)
@@
[…]
@@
$\stackrel{{\scriptstyle 5}}{{=}}\frac{1}{4}\left\{\left[(L^{x})^{T}L^{p}\right]_{ij}^{2}+2(|X_{ij}^{(k)}|-X_{ij}^{(k)})\right\},$ (36)
@@
Combining now Eq. (33) with inequality (4.2.1) we get immediately the following inequality
@@
Finally, putting inequalities (4.2.1) and (40) together, we arrive at the sought necessary condition for biseparability of the form:
@@ 4.2.2 Sum criterion
$\stackrel{{\scriptstyle 2}}{{=}}\sum_{\alpha=x,p}\sum_{l}p_{l}^{(k)}\left[\langle(\Delta u_{i,I_{k}}^{\alpha})^{2}\rangle_{\rho_{I_{k}}^{(l)}}+\langle(\Delta u_{i,J_{k}}^{\alpha})^{2}\rangle_{\rho_{J_{k}}^{(l)}}\right]$
@@
where inequality $1$ again follows from concavity of the variance, whereas in equality $2$ we used the decomposition (34) and the fact that we average over the product state $\rho_{I_{k}\otimes J_{k}}^{(l)}$. To get inequality $3$ we used the inequality between the arithmetic and geometric mean, the uncertainty relations, and the definitions (37) and (38). Finally, if we combine inequalities (44) and (4.2.2) we obtain the desired sum criterion in the following form:
@@
4.3 Direct construction of the sum criterion
@@
4.3.1 Construction of the left-hand side
The quantities $U^{x,p}$, Eq. (20), on the LHS of both criteria are the same and their construction is easy. Namely, each $N$-vertex tree consists of N1N-1 edges [57]. Each of the edges connects a pair of vertices with labels $i$ and $j$, where by $i$ we denote the label which is less than the label $j$, i.e., $i<j$. With each of the edges we associate a two-mode quadrature linear combination $u_{i}^{\alpha}$ (see Fig. 3). Sum of the variances $\langle(\Delta u_{i}^{\alpha})^{2}\rangle$ over all edges then comprises (up to the term $(\ell^{\alpha}_{NN})^{2}\langle(\Delta\alpha_{N})^{2}\rangle$) the quantity $U^{\alpha}$.
@@ 4.3.1 Construction of the left-hand side
The quantity $U^{\alpha}$ in Eq. (20) is then given by the sum of variances of all N1N-1 quantities (50), plus the term $(\ell^{\alpha}_{NN})^{2}\langle(\Delta\alpha_{N})^{2}\rangle$. Making use of this rule we can easily rederive the quantities $U^{\alpha}$ for the linear tree and the star tree given in Eqs. (4.1.1) and (4.1.2), respectively. In the next section we also use the rule to construct the criteria corresponding to the more complex trees in the third column of Tab. 3.
4.3.2 Construction of the right-hand side
@@
Moving to the construction of the RHS it is obvious that to every bipartite split $k\equiv I_{k}|J_{k}$ of $N$ modes into two disjoint subsets $I_{k}=\{i_{1},i_{2},\ldots,i_{l}\}$ and $J_{k}=\{i_{l+1},i_{l+2},\ldots,i_{N}\}$ there corresponds exactly the same bipartite split of the set of all $N$ vertices of the considered graph. The split divides the set of all N1N-1 pairs $(i,e_{ij})$ into two disjoint subsets, denoted as $\mathcal{U}$ and $\mathcal{C}$, depending on whether the edge $e_{ij}$ is uncut by the split or it is cut off by the split. Formally, $(i,e_{ij})\in\mathcal{U}$ if $i,j\in I_{k}$ or $i,j\in J_{k}$, whereas $(i,e_{ij})\in\mathcal{C}$ if $i\in I_{k}$ and $j\in J_{k}$ or vice versa. Now, to each pair $(i,e_{ij})\in\mathcal{U}$ we ascribe the term $|\ell^{x}_{ii}\ell^{p}_{ii}+\ell^{x}_{ji}\ell^{p}_{ji}|$, whereas with each pair $(i,e_{ij})\in\mathcal{C}$ we associate a different term $|\ell^{x}_{ii}\ell^{p}_{ii}|+|\ell^{x}_{ji}\ell^{p}_{ji}|$. The sum of all the latter terms for all N1N-1 pairs $(i,e_{ij})$ plus the term $|\ell^{x}_{NN}\ell^{p}_{NN}|$, which corresponds to the root, gives finally the quantity $\mathcal{K}^{(k)}$, Eq. (47), appearing on the RHS of the sum criterion (46).
@@
5 Applications
5.1 Three-mode states
First, we test functionality of the product criterion (43) and sum criterion (46) for the simplest case of $N=3$ modes. Then there is only one linear tree which for the standard labeling yields the adjacency matrix given above Eq. (4.1.1). Hence, one gets using the formula (41)
$\mathcal{L}=$ $(\ell^{x}_{11}\ell^{p}_{11}+\ell^{x}_{21}\ell^{p}_{21})^{2}+(\ell^{x}_{21}\ell^{p}_{22})^{2}+(\ell^{x}_{22}\ell^{p}_{21})^{2}$
@@ 5.1 Three-mode states
by Eq. (42). Making use of the expressions for the quantities $U_{\rm lin}^{x,p}$, Eq. (4.1.1), we get finally the minimal three-mode product criterion for GME,
$\prod_{\alpha=x,p}$ $\left\{\sum_{i=1}^{2}\langle[\Delta(\ell^{\alpha}_{ii}\alpha_{i}+\ell^{\alpha}_{i+1i}\alpha_{i+1})]^{2}\rangle\right$.
$\quad\left.+\left(\ell^{\alpha}_{33}\right)^{2}\langle(\Delta\alpha_{3})^{2}\rangle\right\}$
@@
$\sum_{\alpha=x,p}$ $\left\{\sum_{i=1}^{2}\langle[\Delta(\ell^{\alpha}_{ii}\alpha_{i}+\ell^{\alpha}_{i+1i}\alpha_{i+1})]^{2}\rangle\right$.
$\quad\left.+\left(\ell^{\alpha}_{33}\right)^{2}\langle(\Delta\alpha_{3})^{2}\rangle\right\}$
$\geq\mathrm{min}\left(|\ell^{x}_{11}\ell^{p}_{11}|+|\ell^{x}_{21}\ell^{p}_{21}|+|\ell^{x}_{22}\ell^{p}_{22}+\ell^{x}_{32}\ell^{p}_{32}|\right$.
$\quad\left.+|\ell^{x}_{33}\ell^{p}_{33}|,|\ell^{x}_{11}\ell^{p}_{11}+\ell^{x}_{21}\ell^{p}_{21}|+|\ell^{x}_{22}\ell^{p}_{22}|\right$.
$\quad\left.+|\ell^{x}_{32}\ell^{p}_{32}|+|\ell^{x}_{33}\ell^{p}_{33}|\right).$ (54)
At the outset we confirm on a particular example that a fully inseparable but not GME state really does not violate the inequalities (5.1) and (5.1). As an example we use the family of fully inseparable biseparable states (3) with CM $\gamma^{\rm test}$, Eq. (12), where $r\in(0,1.24)$. For all states with the squeezing parameter $r_{i}=0.05\times i$, $i=0,1,\ldots,24$, we have minimized numerically the differences $D_{P}$ and $D_{S}$, Eqs. (48) and (49) with $N=3$, i.e., the difference of the LHS and the RHS of inequalities (5.1) and (5.1). The numerical minimization has been done using a built-in function in Wolfram Mathematica [67] called NMinimize where we specified the used optimization method. On a standard desktop computer, the numerical minimization takes from a few seconds in the case of $N=3$ up to $20$ seconds in the case of $N=6$. Those times are approximately the same as for criteria in [36] used on the same devices. Our analysis reveals that for all considered values of the squeezing the differences are very small but always strictly positive numbers, which confirms correctness of the proposed criteria.
@@ 5.1.1 Analytical examples
$\gamma^{\rm sss}=\frac{1}{4}$ $\left[\left(\begin{array}[]{ccc}a_{+}&b_{+}&-b_{+}\\ b_{+}&c_{+}&d_{+}\\ -b_{+}&d_{+}&c_{+}\end{array}\right)\right$. (58)
$\left.\oplus\left(\begin{array}[]{ccc}a_{-}&b_{-}&-b_{-}\\ b_{-}&c_{-}&d_{-}\\ -b_{-}&d_{-}&c_{-}\end{array}\right)\right],$ (62)
@@
Again, we minimized the difference $D_{P}$, Eq. (48), over the parameters $\ell^{\alpha}_{ij}$. To speed up the numerical search we further imposed on the parameters the following constraint $1\leq\ell^{\alpha}_{ij}\leq 1$. The corresponding minimal differences for the CMs (8) and (58) with squeezing parameters $r_{i}=0.05\times i$, $i=0,1,\ldots,40$ are depicted in Fig. 5. The figure as well as the numerics reveal that for the GHZ-like state except for the case $r=0$ for all other considered values of the squeezing the minimized difference is negative. Similarly, in the case of the split squeezed state the minimized difference becomes negative as soon as the considered value of squeezing exceeds the value of $r=0.6$. These two examples clearly demonstrate the ability of the proposed product criterion (5.1) to detect GME of basic three-mode Gaussian states.
@@
Interestingly, the sum criterion (5.1) fails to detect GME of any of the previous three-mode states detected by the product criterion. Nevertheless, one can find examples of three-mode GME Gaussian states, which are detected by both criteria or only by the sum criterion itself.
5.1.2 Numerical examples
We find examples of Gaussian states whose GME can be detected by the sum criterion (5.1) using the Gaussian analog [43] of the iterative algorithm [41]. Initially, a pure-state CM $\gamma^{(0)}$ is randomly generated. In the first step of the algorithm an optimal witness $Z^{(1)}$ detecting GME from minimal set of two-mode marginal CMs of CM $\gamma^{(0)}$ is found. The second step consists of finding an optimal CM $\gamma^{(1)}$ minimizing $\mbox{Tr}(\gamma Z^{(1)})$. In the next iteration the output CM $\gamma^{(1)}$ is taken as an input to the first step which produces an optimal witness $Z^{(2)}$. The new witness is then used in the second step to get optimal CM $\gamma^{(2)}$ for the witness, etc. By performing few iterations of the algorithm on a randomly generated seed CMs we produced one hundred CMs of GME states, which were then tested by our product and sum GME criteria (5.1) and (5.1). In total, our criteria were capable of detecting GME for nearly 80% of the CMs. One of the obtained CMs is given by